1.) Enter your answer in the provided box

Calculate the mass % of O in barium periodate

? % O

2.) Be sure to answer all parts.

Write a balanced equation for the following by inserting the correct coefficients. If the coefficient equals 1, add "1" in the provided box.

??? CaSiO3(s) + ???? HF(g) → ??? CaF2(s) + ??? SiF4(g) + ???H2O(l)

3.) Be sure to answer all parts.

Potassium nitrate decomposes on heating, producing potassium oxide and gaseous nitrogen and oxygen:

4KNO3(s) → 2K2O(s) + 2N2(g) + 5O2(g)

To produce 79.7 kg of O2, how many

(a) moles of KNO3 must be heated?

??mol

(b) grams of KNO3 must be heated?

??g
Answer #1

1.) molecular formula is Ba(IO4)2

molar mass = 519.13 g/mole

mass % of O in barium periodate = (8 * 16) * 100 / 519.13 = 23.66 %

2.) balanced equation is

1 CaSiO3(s) + 6 HF(g) ...............> 1 CaF2(s) + 1 SiF4(g) + 3 H2O(l)

3.) Be sure to answer all parts.

Potassium nitrate decomposes on heating, producing potassium oxide and gaseous nitrogen and oxygen:

4 KNO3(s) → 2 K2O(s) + 2 N2(g) + 5 O2(g)

4 mole KNO3(s) decompose 2 mole K2O(s) and 2 mole N2(g) and 5 mole O2(g).

79.7 kg of O2 = 79.7 * 1000 / 32 = 2490.6 mole.

(a) moles of KNO3 heated = (4 / 5) * 2490.6 = 1.66 * 10^3 mole.

(b) grams of KNO3 must be heated = 1.66 * 10^3 mole * 101.1032 g/mole = 167873 g = 1.68 * 10^5 g